OpenStudy (sweetnikkie072):

Will give medals.. A ferris wheel has a diameter of 10 m and takes 24 sec to make one revolution. The lowest point on the wheel is 1 m above the ground. Sketch a graph to show how a rider's height above the ground varies with time as the ferris wheel makes a rotation. Assume the person starts the ride at the lowest point on the wheel. Write a trigonometric equation that describes the graph.

OpenStudy (anonymous):

\[f(t)=5\cos (\frac{ \pi }{ 12 }t-\pi)+6\]